OpenStudy (anonymous):

A piece of rope was cut into two so that the lengths of the pieces are in the ratio 5:7. If one piece was longer by 18 dm, how long was each piece? How long was the whole rope? explain your answer please :)

OpenStudy (anonymous):

x/y=5/7 7x=5y y=18+x y=18+5/7y y=63 x=81 l=144

OpenStudy (anonymous):

k tnx bro :)

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

y=18+5/7y can you explain this please? i don't get it (i don't know much about math)